OpenStudy (anonymous):

what is the solution of this linear system? 4x+y=-1, y=-2x+3

OpenStudy (texaschic101):

y = -2x + 3 now sub -2x + 3 in for y in the other equation 4x + y = -1 4x + (-2x + 3) = -1 4x - 2x + 3 = -1 -- subtract 3 from both sides 4x - 2x = -1 - 3 -- combine like terms 2x = -4 x = -2 now sub -2 in for x in either equation to find y y = -2x + 3 y = -2(-2) + 3 y = 4 + 3 y = 7 check... 4x + y = -1 4(-2) + 7 = -1 -8 + 7 = -1 -1 = -1 (correct) x = -2 and y = 7 or (-2,7)

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

anytime :)